yomomalikebracobama:

how do i solve 5+3x=100

lilysoul:

3x=95

broidkbutyes:

first you want to eliminate the 5, from the left side of the equation. you subtract 5 and whatever you do on one side you have to do to the other. meaning you have to subtract 5 from 100. 100 - 5 =95. then youre left with 3x=95. you divide 3 on both sides to get x by itself. 95 / 3 = 31.6666 x = 31.6666

wifeofsosaaa:

5 + 3x = 100 \/ subtract 5 from both sides but not the place where x is \/ 3x = 95 \/ then divide 3 by both sides and yu got get one of these answers \/ x = 31.6666666667 or 32

thug4lifebro:

this dude really